# Tileable Surfaces

### v1

@@ 5.2. Prototiles with 3 edges
As mentioned above, there are four different types of interior tile with three edges, depicted in Figure 8. The symmetries of the tiles and the maps that takes a tile to an adjacent tile can be used to study these tilings.
@@
Figure 8. The four interior cases
@@ 6. Conclusions and future directions
In comparison with polyhedral surfaces, smooth tileable surfaces have the attractive feature that they look smooth, but they offer the challenge that they are more difficult to create, due to matching the normals along the edges. We have so far left unexplored finite edge tilings with more than three edges, and tilings that are not finite edge, type as well as larger prototile sets. We expect that more interesting surfaces are obtainable when these are investigated.
@@
    Given an admissible tile, what type of surface(s) does it produce (compact, complete, or incomplete). This is related to the curved surface version of the Heesch problem [16] [21] for plane tilings: given an admissible tile it can be surrounded at least once by a layer of tiles (a corona). If the resulting collection cannot be completely surrounded, then the Heesch number is $1$. In general, the Heesch number is the number of coronas that can be added. Heesch’s tiling problem is: given a positive integer $N$, does there exist a tile with Heesch number $N$? For tileable surfaces we could consider surfaces with or without self-intersections, which would modify the answer for these questions.
@@
    It is known that for tilings of the plane one can find a monotile that generates non-repeating patterns. For curved tileable surfaces, it is easy to give examples of monotiles that generate non-repeating patterns if we consider tiles that are not of finite edge type. Finding finite edge type non-repeating tiles that are not just deformed versions of the planar versions may be an interesting problem.

### v2

@@ 5.2. Prototiles with 3 edges
As mentioned above, there are four different types of interior tile with three edges, depicted in Figure 8. The symmetries of the tiles and the maps that take a tile to an adjacent tile can be used to study these tilings.
@@
Figure 8. The four interior cases.
@@ 6. Conclusions and future directions
In comparison with polyhedral surfaces, smooth tileable surfaces have the attractive feature that they look smooth, but they offer the challenge that they are more difficult to create, due to matching the normals along the edges. We have so far left unexplored finite edge tilings with more than three edges, and tilings that are not finite edge type, as well as larger prototile sets. We expect that more interesting surfaces are obtainable when these are investigated.
@@
    Given an admissible tile, what type of surface(s) does it produce (compact, complete, or incomplete). This is related to the curved surface version of the Heesch problem [heesch1968regulares] [mann2004heesch] for plane tilings: given an admissible tile it can be surrounded at least once by a layer of tiles (a corona). If the resulting collection cannot be completely surrounded, then the Heesch number is $1$. In general, the Heesch number is the number of coronas that can be added. Heesch’s tiling problem is: given a positive integer $N$, does there exist a tile with Heesch number $N$? For tileable surfaces we could consider surfaces with or without self-intersections, which would modify the answer for these questions.
@@
    It is known [smith2023aperiodic] that for tilings of the plane one can find a monotile that generates non-repeating patterns. For curved tileable surfaces, it is easy to give examples of monotiles that generate non-repeating patterns if we consider tiles that are not of finite edge type. Finding finite edge type non-repeating tiles that are not just deformed versions of the planar versions may be an interesting problem.